Server and Network Status (page refreshes every 60 seconds) We may also report the status on: Follow seirettostatus on Twitter

Seiretto client area and renewals - planned maintenance:

       There is currently no planned maintenance.

Designer/reseller/DNS hosting accounts:

Current System Status

       All systems are operating normally.

Planned System Maintenance

       There is currently no planned maintenance.

 Hybrid & VPS servers, Premier, Small business and Starter accounts:

Current System Status

All systems are operating normally.

Planned System Maintenance

There is currently no planned maintenance.


Recent resolved issues/maintenance

14:46 24/02/2017 - E-Mail Delivery to BT and Yahoo Identified - We are seeing that a number of rejections to BT and Yahoo are being caused by badly formed SPF records. If you are having issues sending to addresses with BT and Yahoo from our systems we recommend using the following SPF:-

v=spf1 exists:%{i} exists:%{i} -all

If you need any assistance in updating this our support team will be happy to help.  

22:20 09/02/2017 - Hotmail/Outlook and other Microsoft addresses sending issues - affecting some Premier, Small business and Starter accounts - This is still ongoing, we are continuously monitoring this and will keep this page up to date with the latest information. Investigating - We are currently seeing a large amount of emails being deferred when sending to Hotmail/Outlook and other Microsoft addresses. This includes and domains that have emails hosted with outlook services. This looks to be caused by a limit on their servers restricting the amount of emails that can be received from our shared mailservers. Our sytem admin team are currently investigating and attempting to contact them for further details, or a possible work around. On inspection this may not be effecting all emails sent to these addresses, but a large amount are being blocked. This will cause bouncebacks to the sender, so you can see which emails have been blocked. Our apologies for any inconvenience caused here in the meantime.

5.24pm, 27/01/2017 We are currently seeing abnormal behaviour on VPS machines. We expect to have it back to normal operation soon, and apologise for any inconvenience caused. Update -> We have identified that two disks have failed on KVMhost 69, we are currently look at restoring the data however due to the size of the data we have a rough estimate of 24 hours for this to be completed. Update 10:07 -> We are currently still restoring the data, we will provide you further updates shortly, we apologise for any inconvenience caused regarding this.  Recovery is ongoing but the success rate is currently around 20% and some VPSs are being restored from backups.

1pm 17/18th Jan 2017
- Connectivity Issues - affecting Premier/small business/starter accounts, rebuilt being applied.  Some NAS Drive synchronisation is still taking place, this will cause high load on some servers which may cause intermittent problems accessing some sites. Any access issues should only be momentary

Update to client area and renewals
7pm  6th Jan 2017
- Update to our client area: and renewals:
It is expected to take 1 hour.  We apologise for any inconvenience caused during this time.

9.38am 12th Dec 2016 - Domain Renewals, Registrations and Updates

We are currently seeing delays with some domain registrations, renewals and contact changes. Our Systems Administration Team are working on this issue and we expect to have this fully corrected shortly.
Your patience is appreciated and we apologise for the inconvenience caused.

1:00 GMT, Thursday, 8th December, 2016 - Migration.
We will be migrating all Designer accounts to new faster servers. As the migration will start at 1am we expect the whole process to complete by normal business hours (8-9am).
You should ensure you backup now, or the day before the migration, so you have a recent copy of your data, please see: 
No other action is required by you. However, during the migration there will be a period where the DNS is switching across to the new servers the requests can go from visitors to either server, so updates to sites, etc. should not be made during this time and if you have an ecommerce store orders may come through to either one (if this could be an issue for your website then it may be best to treat it as a maintenance window and not take orders during this time).

7am 8th Dec -  all accounts now copied to new servers, other than minor issues all systems are working as expected. 
Any domains not using our nameservers ( - must update the DNS A records to point to (replacing or

8pm Wedneday 7th December to 8am Thursday 8th December 2016
We will be updating our members area and renewals backend systems.   During this period you will be unable to login or complete renewals, please return when updates are complete.  Apologies for any inconvenience this causes.

22nd Aug 2016 - scheduled upgrades

starting at 8pm we will be performing scheduled upgrades to our shared Linux hosting environment, server hardware will be upgraded so websites may experience slowness/brief periods of downtime. We apologies for any inconvenience caused.

3pm, 10th Feb 2016 - Power outage

There was a brief interruption in power during works in one of our data centre halls. Which is affecting VPS servers, Premier, Small Business and Starter accounts. 
This has now been restored and all servers are coming back online.  However, ramifications continued for several hours as servers rebooted and multiple errors occured.   11.50am 11th Feb: Service is returning to normal slowly.

11th Jan 2016 - Traffic from AWS Network

Over the Christmas period we experienced an extremely large amount of traffic in the form of a DOS attack which has originated from AWS's network. As such, to protect our services, we have had to block all traffic from that network.  Our systems admin team have contacted the AWS abuse team to try and resolve this issue. Once the AWS abuse team have been able to identify the origin within their network for the attack, and resolve the issue, we can then safely enable traffic from their network again.   Apologies for any inconvenience this causes.

9pm 28th Oct 2015 - Planned System Maintenance mail174 mail179
We will be rebooting these mail servers at the above time to perform routine maintenance. Mail services will be unavailable during this time. We expect the work to take no longer than an hour.

14th Oct 2015, 10pm - Shared Windows servers
 At the the above time and date the Windows Shared Web servers will be rebooted in order to apply critical updates. This should take around 30 minutes. We apologise for any inconvenience.

3rd Aug 2015 -Windows "web211" - Abnormal Server Behaviour

We are currently seeing abnormal behaviour on this server, and are investigating the causes.  We expect to have it back to normal operation soon, and apologise for any inconvenience caused.

*Update* 2015-08-01 17:20  We are continuing to investigate the issue, we apologise for the inconvenience caused and hope to have this running as normal again shortly.

*Update* 2015-08-02 11:10  We are still continuing to investigate the issue, as a top priority, and hope to have this server back in service as soon as possible.

*Update* 2015-08-03 08:08  Work is still continuing on this issue, we have our most senior engineers working on this as a top priority. We will have a more detailed and official report before 10am here. We sincerely apologise for the inconvenience caused.

10pm 13th July 2015 -  Planned System Maintenance MSSQL2 Reboot - will affect some Small business and Premier accounts.
At the above time and date this server will be rebooted in order to apply important updates. We expect this to take around 30 minutes.

25th June 2015 - abnormal servers performance - some Starter, Small business and Premier accounts affected (web150, web151, web152)
We are investigating the causes, normal operation expected soon.  We apologise for any inconvenience caused.

Update: 25/06/2015 (16:25) - We have suffered a critical failure for this cluster which has resulted in a corrupt file system. During this time access to websites will be intermittent. We're working to fix this and restore access as quickly as possible. Please accept our apologies for any inconvenience caused, and please be assured this incident is being treated with our highest priority.

Update: 25/06/2015 (22:43) - Our system administrators are still working on this with the highest priority and we endeavour to try and resolve this as quickly as possible. Once again, apologies for any inconvenience this has undoubtedly caused.

Update: 26/06/2015 (08:00) - The work is still ongoing and at present a large number of customer sites are back online.

Update: 26/06/2015 (9:50) - Websites are currently being restored on an alphabetic basis. We expect this to complete no later than the 27th June, our estimated time of completion being 16:00-18:00.

10th June 2015 - Windows Shared Hosting updates
At 10pm the shared windows servers will be rebooted to apply critical updates, this will take around 1 hour.

10th April 2015 - affecting some Starter, Small Business and Premier accounts:

web175, web176, web177, web178 abnormal servers behaviour, currently seeing abnormal behaviour on these servers. Websites that do not require making updates to files in the hosting space to load will be online however customers will be unable to make changes in their hosting space via FTP or  Filemanager.

28th March 2015 - update to our client area:
Work will start sometime after 3pm and is expected to take 2 hours.


^back to top^

Copyright © 1996-2017 Seiretto Ltd. All rights reserved.
Registered in England & Wales no: 4716409. VAT no: GB780 4245 32